OpenStudy (anonymous):

True or False: The point (0, 7) is a solution to y ≤ -3x + 7.

OpenStudy (anonymous):

If you put the x=0 and y=7 values into the inequality, you will have: \[7 \le 3*0+7\] \[7 \le 7\]

OpenStudy (anonymous):

Sorry, that was meant to be -3*0, but the result stays the same.

OpenStudy (anonymous):

We have \[y=7\] And : \[-3x+7=-3\times0+7=7\] And because \[7\leq7\] the (0,7) is a solution to the inequality !

OpenStudy (anonymous):

so it is true